[figure: 9.7 cm]

Normal state spin dynamics in Sr_{2}RuO_{4}: anisotropy due to spin-orbit coupling

I. Eremin, D. Manske, and K.H. Bennemann

Institut für Theoretische Physik, Freie Universität Berlin, D-14195 Berlin, Germany

The novel spin-triplet superconductivity with T_{c}=1.5K observed in layered Sr_{2}RuO_{4} seems to be a new example of unconventional superconductivity. Most recently it was found that the superconducting order parameter contains line nodes half-way between the RuO_{2}-planes[1]. These results support a scenario of Cooper-pairing via spinfluctuations as one of the most probable mechanism to explain the superconductivity in Sr_{2}RuO_{4}. Using the three-band Hubbard Hamiltonian we calculate within the random-phase-approximation the spin susceptibility, χ(q, ω) and NMR spin-lattice relaxation rate, 1/T_{1}, in the normal state of Sr_{2}RuO_{4} and find quantitative agreement with experimental data[2]. Most importantly, at low temperatures we find that the out-of-plane component of the spin susceptibility, χ^{zz} [3], becomes two times larger than the in-plane one, χ^{+−}/2, due to spin-orbit coupling. Furthermore, we obtain that the spin-orbit coupling induced anisotropy results in incommensurate antiferromagnetic fluctuations pointing into the z-direction originated from the quasi-one-dimensional xz- and yzbands. Our results provide further evidence for the importance of spin fluctuations for triplet superconductivity in Sr_{2}RuO_{4}.

[1] M.A. Tanatar et al., Phys. Rev. Lett. 86, 2649 (2001).

[2] K. Ishida et al., Phys. Rev. B 64,100501(R) (2001).

[3] I. Eremin, D. Manske, and K.H. Bennemann, preprint cond-mat/0102051.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : I. Eremin

Address for correspondence : Institut für Theoretische Physik Freie Universität Berlin Arnimallee 14, D-14195 Berlin Germany

Email address : ieremin@physik.fu-berlin.de